Solve:
x squared + 11x = -28

Answer by jim_thompson5910(35256)

x^2 + 11x = -28

x^2 + 11x + 28 = 0

(x+4)(x+7) = 0

x+4 = 0 or x+7 = 0

x = -4 or x = -7